# Tree networks of real-world data: analysis of efficiency and spatiotemporal scales

### v4

@@ Abstract
This work computationally investigates an algorithm that was originally proposed for creating and searching trees of patterns under the assumption that it seeks the most efficient internal representation of the external world, namely a real-world dataset of patterns. The algorithm produces a transition from star graph to regular 1d lattice, passing through irregular scale-invariant trees, under the effect of a single control parameter. The relatively novel information-theoretic tools of the von Neumann entropy and its derivative, the specific heat, together with the recent definition of thermodynamic quantities, such as work and heat, for networks are reviewed and applied to inspect the tree structures produced by the formation protocol. We observe some consistency between the optimal thermodynamic efficiency for networks and the optimal pattern-matching efficiency, defined in terms of the trade-off between accuracy and computational cost when the former is measured at the meso-resolution scale involved in the phase transition from star graph to scale-invariant network. We adduce the hypothesis of a richer frequency content at the meso-scale of the self-organized network to explain this observation. Moreover, empirical evidence suggests that the system benefits from the vicinity to the mentioned phase transition if it is faced with higher-complexity environments, while it does not when the environment is simpler. A phase portrait of the tree-generating algorithm is provided and multiple phase transitions are identified.
@@
Keywords Trees  $\cdot$ Networks  $\cdot$ Thermodynamics  $\cdot$ Statistical physics  $\cdot$ Machine Learning  $\cdot$ von Neumann entropy  $\cdot$ Specific heat
1 Introduction
@@
Why complex structures emerge in a real-world environment is a fascinating question that has not yet found a fully satisfactory answer. In a recent article [1], the emergence of complex structures is devised as the macroscopic thermodynamic result of the optimal ability to store the information about the environment in an accessible manner with the ability to transmit information across the structure [2]. The latter is defined in terms of microscopic quantities thanks to the fairly new developments in statistical physics in network theory [1, 3, 4, 5]. In living systems, natural selection is the way the information about the environment flows into the system’s genome which will then encode such information in the structural traits of the organisms which in turn reflect the functional capability of the system. Thus, the structural traits of a system could themselves enclose the complexity of the environmental conditions the system needs to adapt to, or in other words, to represent into its very own structure [6]. How these traits are optimized to maximize the fitness to the environment is an open question. The hypothesis of optimal functioning at the border of order and disorder suggests the critical point as the potential candidate to obtain functional advantages [7, 8, 9, 10, 11] and maximize the fitness to the outer world, or to obtain a better performance in a statistical generalization problem in the context of learning machines. Here we challenge this hypothesis. In our work, a system seeks the best-fitness configuration given some external constraints $r$ and a high-dimensional vector collection, the dataset $\mathcal{D}$.
We take as a case study the algorithm introduced by Portegys [12] for accelerating nearest-neighbor (1-NN) pattern recognition tasks and subsequently extended to k-NN [13]. The algorithm employs a tree-formation protocol where raw data are reorganized into hierarchical tree structures using a recursive tree-formation method in which samples are randomly selected and inserted into the growing network. At each step, the new pattern is linked to either the current parent pattern or to one of the k-NN children of the parent pattern. The choice is based on the distance between the parent and the new pattern and the one evaluated between the child pattern and the new pattern. The parameter $r$ modulates the relative importance of these two distances creating a larger or smaller region of influence of the parent in the sample space. Figure 1 shows the three structure prototypes produced depending on the value of the control parameter $r$. For $r\sim 0$, every node is linked to the root node (see the central node, which is the first node selected in the process) and a star graph turns out. For some intermediate value of $r$ a ragged tree structure with progressively smaller distances between the patterns is produced, while for $r\rightarrow 2$ a 1d-lattice is created. The algorithm produces exclusively trees, i.e. loop-less networks, which are necessarily sparse $|E|=N-1$ where $|E|$ and $N$ are the number of links and nodes respectively. Using the metaphor introduced by Jacob for natural selection [14], the algorithm does act as a tinkerer that creates a structure using what is available at the time, collocating one node after the other across the tree formation process according to a parameterized distance-based rule.
@@
The use of a hierarchical topology, onto which a high-dimensional dataset is mapped, offers insights into the structured disorder of the data [15] on the one hand, while on a practical perspective, it limits the number of comparisons needed in matching an unknown pattern to a prototypical example. Namely, the algorithm introduced by Portegys is equipped with a pattern-searching method described in Section 3.1. Depending on the structural traits of the networks, an optimal trade-off between fast search of the best-matching pattern and accuracy in terms of node-class labels can be defined.
The Portegys algorithm has the desirable properties of being simple, explainable, and subject to a single control parameter $r$. Thanks to these features, we can work on the bridge between three disciplines: statistical physics, machine learning, and network science, and we can show that information-theoretical tools such as the von Neumann entropy and its derivative reflect the functional capabilities of the network. Notably, we observe some consistency between the thermodynamic efficiency and pattern-matching efficiency that we try to explain in terms of average frequency content in the self-organized network structure.
2 Theoretical framework
2.1 Statistical physics of networks
@@
As $\hat{K}$ is a positive semidefinite Hermitian matrix admits the spectral decomposition in terms of the sum of projectors $\hat{K}=\sum_{i}^{N}e^{-\tau\lambda_{i}}|k_{i}><k_{i}|$. Thus, for any state at $\tau$ it holds $|\phi(\tau)>=\sum_{i}^{N}e^{-\tau\lambda_{i}}|k_{i}><k_{i}||\phi(0)>$. This relation remarks that the eigenvectors of the Laplacian are akin to the wave vectors in the Fourier space defined by the spectrum of the continuum-Laplacian, while $\tau$ can be interpreted as the "wavelength" $l\sim\tau$ of the diffusion modes, thus it defines an operational resolution scale for observations [3]. Moreover, the ensemble of accessible diffusion modes at time $\tau$ can be constructed by defining a probability measure over the accessible modes decomposed by the spectral theorem. Such canonical ensemble [5, 4, 22] is encoded into the density matrix:
@@
The elements $\rho_{ij}$ represent the normalized amount of information transferred in a diffusion process between nodes $i$ and $j$ at time $\tau$. Remarkably $\rho_{ij}$ takes into account all possible topological pathways between the two nodes and assigns smaller weights to longer ones tantamount to path-integrals [20, 22]. The partition function $Z=\operatorname{Tr}[\hat{K}]=\sum^{N}_{i=1}e^{-\tau\lambda_{i}(\hat{L})}$ is a function of the eigenvalues, $\lambda_{i}$, of the Laplacian and is proportional to the average return probability of a continuous time edge-centric random walker to be back in its initial location after time $\tau$ [4]. Note that the object $\hat{\rho}$ is a positive semi-definite and Hermitian matrix whose trace sums to unity which has allowed its eigenvalues to be interpreted as probabilities.
@@
where $\{\mu_{i}(\tau)\}_{i=1}^{N}$ is the set of $\hat{\rho}(\tau)$ eigenvalues that are related to the set of Laplacian eigenvalues through $\mu_{i}(\tau)=e^{-\tau\lambda_{i}}/Z$. The entropy $S(\hat{\rho}(\tau))$ is a function of the normalized time $\tau$ and is bounded between $[\log C_{conn},~{}\log N]$ where $C_{conn}$ is the number of connected components and $N$ is the number of nodes. Thus the von Neumann entropy, $S$, as a function of $\tau$ reflects the entropic phase transition of information propagation (or loss) over the network [22]. Specifically for a connected network, at $\tau\rightarrow 0$, $S(\tau)=\log N$ reflects the segregated heterogeneous phase where information diffuses from the single nodes to the very local neighborhood; at $\tau\rightarrow\infty$, $S(\tau)=0$ the diffusion is governed by the smallest non-zero eigenvalue of the Laplacian, associated with the so-called Fiedler eigenvector, reflecting the integrated homogeneous phase where the information has propagated all over the network.
@@
The network entropy thus counts the (logarithmic) number of informationally unconnected nodes in the heterogeneous network and by the inspection of the entropy $S$ and its generalized susceptibility (or specific heat), $C=-dS/d\log\tau$, it is possible to identify the characteristic resolution scales of the network [22, 3]. For instance, a single large peak of the specific heat $C$ identifies a single scale where there is a fast and unique information diffusion (or information lost) across the network. The latter is the typical situation of a star graph, depicted in Figure 1, or an Erdös-Rényi network. Before the peak, the informationally connected components are the single nodes, while after the peak the informationally connected components are the full network revealing a single trivial scale. On the contrary, a constant specific heat reflects the scale-invariance of the network in a resolution region where information is lost at a constant rate. Such a situation is typical of homogeneous $d$-dimensional lattices and of heterogeneous networks whose spectral density is power-law distributed in a sufficiently extended scale-interval [3]. The latter sentence makes for a definition for the scale-invariant networks [23]. The specific heat plateau height, $C_{0}$, is also related to the spectral dimension of the network $d_{s}=2C_{0}$ [3, 23].
@@
From these considerations, a Laplacian-Renormalization group theory has been recently developed allowing for a coarse-graining of the network structures into super-nodes unraveling the mesoscale structures [3, 22]. In this context, the peaks of the specific heat, $C$, or a marked change in its slope, identify the most convenient resolution scales where to perform the coarse-graining over the informationally connected nodes, as it allows the iteration of the procedure without an excessive network reduction, and, when the network is scale-invariant, maintaining the dilation symmetry across the renormalization group procedure [19].
@@
To conclude, we can intuitively use the diffusion parameter $\tau$ to resolve the network topological structures at different resolution scales. Thus, the necessity of analyzing complex networked systems at all levels is solved by resorting to the conjugate Fourier space, the k-space, defined by the Laplacian spectrum and the "wavelengths" of its modes.
2.2 Thermodynamics of networks
@@
The statistical physics theory for networks introduced above allows us to view network phenomena from a new perspective and possibly to bridge the gap to network macroscopic observables. To this regard network thermodynamic quantities such as work, $W$, and heat, $Q$, have been defined in mathematical analogy to their thermodynamic counterparts [1]. Since a network of isolated nodes is the only network with maximum entropy $S_{iso}=\log N$ for $\tau>0$, progressively adding edges diminishes the network’s entropy. Where lower entropy reflects the diminished ability of the network to express diversity to a stochastic perturbation, or in other terms, the loss in the internal representation capabilities encoded in the network topological traits. The entropic cost is given by:
@@
where $0\leq\eta\leq 1$ and $U\equiv-\partial_{\tau}\log Z\equiv\operatorname{Tr}[\hat{L}\hat{\rho}]$. The variational principle here exposed has been devised to macroscopically explain the sparsity and its empirical scaling law, as well as the emergence of topological features such as modularity, small-worldness, and, with partially less clear results, heterogeneity [1]. In our work, we use this variational principle to explain the structural properties of the network subclass composed of heterogeneous hierarchical trees. An additional element is that these trees are required to encode the diversity expressed in the environment they are in contact with, defined by a real-world dataset of patterns. The specific heat is used here to identify which is the resolution scale at which such a variational principle is effective for our subclass of networks during the transition to the scale-invariant regime. For scale-invariant networks, for some $\tau$-scales, $U=<\lambda>_{\tau}=C_{0}/\tau$ [3, 23], hence in this regime Equation 5 becomes $\eta=\frac{1}{\tau}\frac{C_{0}}{W}$.
3 Methods
Our results are evaluated over subsets of four different datasets: the NIST [25], the MNIST [26], the FashionMNIST [27], and the CIFAR10 [28] datasets. These datasets contain randomly collected instances of image samples out of 10 classes. It should be noted that in this type of data, a pattern class can generally not be represented by a single centroid vector, but will contain multiple densities (’sub-style patterns’) such that the sample point cloud can be ordered naturally in a network structure.
Figure 2: Dataset complexity is estimated with multiple methods for each training dataset of 2000 samples encountered in this work. In green an estimate is given computing the Lempel-Ziv complexity [29], while in orange the intrinsic dimension is estimated with the Maximum Likelihood Estimation approach [30, 31, 32]. In blue the accuracy on the test set for a multilayer-perceptron (MLP) with two hidden layers of 128 units each. The left axis refers to the quantities in green and orange, while the right axis refers to the multilayer-perceptron accuracy. We observe that the CIFAR10 dataset is the most complex one, while the NIST dataset is the least complex among the four.
The MNIST database contains $24\times 24$ grey-scale images of handwritten digits [0-9] commonly used to benchmark machine- and deep-learning algorithms. The NIST dataset is the ancestor of the MNIST, it contains $16\times 16$ black-and-white images of the handwritten digits [0-9]. The FashionMNIST is a large dataset of $28\times 28$ gray-scale images of fashion products from 10 categories. The CIFAR-10 dataset contains $32\times 32$ color images belonging to 10 classes. In all our experiments a subset of 2000 samples is extracted from the train set and the test set independently. Samples from the FashionMNIST and MNIST are always binarized, while the samples from the CIFAR10 are normalized between 0 and 1 dividing each pixel value by 255, and values are rounded up to eighth decimal. In the Supplementary we show also the same results for larger networks generated from $4000$ samples from the MNIST and FashionMNIST. The 10 classes are equally represented both in the test and in the training set. The datasets used in our experiments exhibit different levels of complexity, see Figure 2. We estimate with multiple methods the dataset complexities on the same subset of 2000 training samples used in the rest of the manuscript. The first heuristic approach is based on the test accuracy of a small multilayer perceptron (2 layers with 128 units each). In pattern recognition, the complexity of a dataset is determined by the complexity of the curvilinear hyperplanes that would be needed to optimally separate patterns from different classes. A simple but effective parameter to gauge this facet of complexity is the accuracy that a classifier (such as the MLP) can achieve in determining the true class label for a data point. Thus, a more complex dataset would result in a lower accuracy for neural networks of fixed size when compared to a less complex dataset. The second method we use is the computation of the Lempel-Ziv complexity [29] of the aggregated training set. In this case, samples need to be binarized. The third method concerns the measure of the intrinsic dimension [15, 32] of the dataset using the Maximum Likelihood Estimation approach introduced in [30] with the correction proposed in [31]. A given dataset of gray-scale images of size $pixel\times pixel$ a priori lies on a high-dimensional space of $\mathbb{R}^{pix\times pix}$, nevertheless, because it might contain a specific class of objects that account for the "world" of our problem, for instance, handwritten digits, it spans a lower dimensional manifold. Even though the definition of such a manifold is an open problem, it is possible to estimate its dimension based on the distance of the $n$-neighbouring points [15, 32]. We set the hyper-parameter $n=3$.
While in Section 4 we discuss the full range of $r$, in Section 5 we limit the discussion to $r\leq 0.85$ as the accuracy, and hence the pattern-matching efficiency, quickly drops for larger $r$, moreover the algorithm shows instability as it doesn’t necessarily converge for $r\geq 0.9$. The Portegys algorithm we use is the extended version with $k=5$ introduced in [13] where the distance between patterns is the Euclidean distance.

### v5

@@ Abstract
Hierarchical tree structures are common in many real-world systems, from tree roots and branches to neuronal dendrites and biologically inspired artificial neural networks, as well as in technological networks for organizing and searching complex datasets of high-dimensional patterns. Within the class of hierarchical self-organized systems, we investigate the interplay of structure and function, associated with the emergence of complex tree structures in disordered environments. Using an algorithm that creates and searches trees of real-world patterns, our work stands at the intersection of statistical physics, machine learning, and network theory. We resolve the network properties over multiple phase transitions and across a continuity of scales, using the von Neumann entropy, its generalized susceptibility, and the recent definition of thermodynamic-like quantities, such as work, heat, and efficiency. We show that scale-invariance, i.e. power-law Laplacian spectral density, is a key feature to construct trees capable of combining fast information flow and sufficiently rich internal representation of information, enabling the system to achieve its functional task efficiently. Moreover, the complexity of the environmental conditions the system has adapted to is encoded in the value of the exponent of the power-law spectral density, inherently related to the network spectral dimension, and directly influencing the traits of those functionally efficient networks. Thereby, we provide a novel metric to estimate the complexity of high-dimensional datasets.
@@
Trees, Networks, Thermodynamics, Statistical physics, Machine Learning, von Neumann entropy, Specific heat
I Introduction
@@
Why complex structures emerge in a real-world environment is a fascinating question that has not yet found a fully definite answer. In a recent article [1], the emergence of complex structures is devised as the macroscopic thermodynamic result of, on the one side, the optimal ability to store information about the environment in an accessible manner, on the other side, the ability to transmit information across the structure [2] (see Methods Sec. II.1.2). For instance, in living systems, information about the environment flows into the system’s genome through natural selection. The genome encodes such information into the organism’s structural traits, which in turn reflect the functional capability of the system. Thus, the system’s very own structure may be regarded as the internal representation of the environment [3, 4, 5, 6, 7] and its the structural traits could themselves enclose the complexity of the environmental conditions the system needs to adapt to. How these traits are optimized to maximize the fitness to the environment is an open question. The hypothesis of optimal functioning at the border of order and disorder suggests critical-like features as the potential candidates to obtain functional advantages [8, 9, 10, 11, 12, 13, 14] and maximize the fitness to the outer world, or, when in the context of learning machines, to obtain a better performance in a statistical generalization problem [10]. Among the various attributes exhibited by truly critical systems and systems with critical-like properties, especially relevant to our work is the property of scale-invariance, exhibited through the power-law distribution of some meaningful quantity of the system under consideration.
@@
Hierarchical trees enable us to investigate the interplay between structure and function, associated with the emergence of complex structures in disordered environments, within a relevant class of systems. Trees are common to many real-world systems, from tree roots and branches, to neuronal dendrites [15, 16, 17, 18, 19] and biologically inspired artificial neural networks [20, 21, 22, 23, 24, 25], as well as technological networks for organizing and searching collections of high-dimensional patterns [26, 27, 28].
We use the Portegys algorithm [27, 28] to produce hierarchical trees that self-organize given a collection of high-dimensional vectors. The algorithm recursively organizes (structural aspect) high-dimensional vectors randomly drawn from a dataset into trees (see in Methods Sec. II.2). Using the metaphor introduced by Jacob for natural selection [29], the Portegys tree-generation process acts as a tinkerer that creates a structure using what is available at the time, collocating one node after the other across the tree formation process according to a parameterized distance-based rule. The algorithm integrates a search phase (functional aspect) where it navigates the tree structure to achieve efficient pattern-matching.
The algorithm is simple, explainable, and subject to a single control parameter, $r$. Still, it is capable of generating a large variety of branching structures (Fig. 1) and multiple phase transitions (Sec. III.3). Other models can generate hierarchies. Such as the model in Ref. [30] where a single parameter controls the branching and produces structures in much analogy to the ones in Fig. 1, or the Galton–Watson stochastic process originally proposed to model the extinction of family names  [31, 32]. Nevertheless, the Portegys algorithm integrates the functional aspect given by the pattern-matching task and includes the disorder of the external dataset, two key attributes in our analyses. Thanks to these features, we can work at the intersection of three disciplines: statistical physics, machine learning, and network science. We can show that information-theoretical tools such as the von Neumann entropy [33] and its generalized susceptibility [34, 35] (see Methods), while monitoring the structural properties of the system, reflect the functional capabilities of the network.
We show that emerging efficient structures exhibit a scale-invariant spectrum [36], i.e. power-law Laplacian spectral density. We reproduce this result over multiple datasets, ranked based on their complexity. Our findings suggest that a large range of availability in the power law spectral density provides the advantage of a large repertoire of network “frequencies”. These can be exploited to achieve the functional purpose of the system efficiently, akin to critical-like systems exhibiting scale-invariance in the distribution of a relevant quantity [37, 11, 38, 13, 14]. Such behavior is observed consistently whether we take efficiency as gauged by the task-agnostic network thermodynamic efficiency [1], estimating the trade-off between the diversity of information pathways and fast information propagation (see Sec. II.1.2 in Methods), or the pattern-matching efficiency (Sec. II.2.2 in Methods), quantifying the trade-off between accuracy and number of visited nodes within the pattern-matching task.
Furthermore, the exponent of the power-law spectral density, related to the network generalized dimension [39, 40, 41, 42], results to be driven by the complexity of the environment the system adapted to, directly influencing the traits of those functionally efficient networks.
Finally, from a different perspective, using a hierarchical topology to map the distribution of high-dimensional vectors of a dataset provides insights into the structured disorder of the data [43]. In particular, we measure the dataset network’s spectral dimension and show that it estimates the dataset’s complexity. Thus, among the other results, our work provides an alternative way to gauge the dataset’s complexity [43, 44, 45, 46, 47, 48].
II Methods
II.1 Theoretical framework
II.1.1 Statistical physics of networks
@@
As $\hat{K}$ is a positive semidefinite Hermitian matrix admits the spectral decomposition in terms of the sum of projectors $\hat{K}=\sum_{i}^{N}e^{-\tau\lambda_{i}}|k_{i}><k_{i}|$. Thus, for any state at $\tau$ it holds $|\phi(\tau)>=\sum_{i}^{N}e^{-\tau\lambda_{i}}|k_{i}><k_{i}||\phi(0)>$. This relation remarks that the eigenvectors of the Laplacian are akin to the wave vectors in the Fourier space defined by the spectrum of the continuum Laplacian. At the same time, $\tau$ can be interpreted as the “wavelength”, $\ell\sim\tau$, of the diffusion modes, thus it defines an operational resolution scale for observations [35]. Moreover, the ensemble of accessible diffusion modes at time $\tau$ can be constructed by defining a probability measure over the accessible modes decomposed by the spectral theorem. Such canonical ensemble [33, 54, 34] is encoded into the density matrix:
@@
The elements $\rho_{ij}$ represent the normalized amount of information transferred in a diffusion process between nodes $i$ and $j$ at time $\tau$. Remarkably $\rho_{ij}$ takes into account all possible topological pathways between the two nodes and assigns smaller weights to longer ones tantamount to path-integrals [52, 34]. The partition function $Z=\operatorname{Tr}[\hat{K}]=\sum^{N}_{i=1}e^{-\tau\lambda_{i}(\hat{L})}$ is a function of the eigenvalues, $\lambda_{i}$, of the Laplacian and is proportional to the average return probability of a continuous time edge-centric random walker to be back in its initial location after time $\tau$ [54]. Note that the object $\hat{\rho}$ is a positive semi-definite and Hermitian matrix whose trace sums to unity allowing its eigenvalues to be interpreted as probabilities.
@@
where $\{\mu_{i}(\tau)\}_{i=1}^{N}$ is the set of $\hat{\rho}(\tau)$ eigenvalues that are related to the set of Laplacian eigenvalues through $\mu_{i}(\tau)=e^{-\tau\lambda_{i}}/Z$. The entropy $S(\hat{\rho}(\tau))$ is a function of the normalized time, $\tau$, and is bounded between $[\log C_{conn},\leavevmode\nobreak\ \log N]$ where $C_{conn}$ is the number of connected components and $N$ is the number of nodes. Thus the von Neumann entropy, $S$, as a function of $\tau$ reflects the entropic phase transition of information propagation (or loss) over the network [34]. Specifically for a connected network, at $\tau\rightarrow 0$, $S(\tau)=\log N$ reflects the segregated heterogeneous phase where information diffuses from the single nodes to the very local neighborhood; at $\tau\rightarrow\infty$, $S(\tau)=0$ and the diffusion is governed by the smallest non-zero eigenvalue of the Laplacian, associated with the so-called Fiedler eigenvector, reflecting the integrated homogeneous phase where the information has propagated all over the network.
@@
The network entropy thus counts the (logarithmic) number of informationally unconnected nodes in the heterogeneous network. By the inspection of the entropy, $S$, and its generalized susceptibility (or specific heat), $C=-dS/d\log\tau$, it is possible to identify the characteristic resolution scales of the network [34, 35]. For instance, a single large peak of the specific heat, $C$, identifies a single scale at which there is a fast and unique information diffusion (or information loss) across the network. The latter is the typical situation of a star graph, depicted in Fig. 1, or an Erdös-Rényi network. Before the peak, the informationally connected components are the single nodes, while after the peak the informationally connected components are the whole network, thus revealing a single trivial scale. On the contrary, a constant specific heat reflects the scale-invariance of the network in a resolution region where information is lost at a constant rate. Such a situation is typical of homogeneous $d$-dimensional lattices and of heterogeneous networks whose spectral density is power-law distributed, $\mathcal{P}(\lambda)\sim\lambda^{\gamma}$, in a sufficiently extended scale-interval [35]. The latter sentence makes for a definition for the scale-invariant networks [36], characterized by a scale-invariant information propagation across the network resolution scales. The the specific heat plateau height, $C_{0}$, is also related to the spectral dimension, $d_{s}=2(\gamma+1)$ [39, 41], according to the relation $d_{s}=2C_{0}$ [35, 36], which implies $C_{0}=\gamma+1$. The spectral dimension generalizes the concept of the Euclidean dimension to disordered structures and coincides with the Euclidean dimension for lattices. For instance, we expect $C_{0}=1/2$ for 1d-lattice; $C_{0}=1$ for the regular grid with periodic boundary conditions; and $C_{0}=4/6$ for random trees with finite degree variance $<k^{2}>$ showcasing fractal dimension $d_{s}=4/3$ [36].
@@
From these considerations, a Laplacian-Renormalization group theory has been recently developed, allowing for a coarse-graining of the network structures into super-nodes and for the unraveling of the mesoscale structures [35, 34]. The theory relies on information diffusion to identify informationally equivalent “blocks” akin to the ones proposed by Kadanoff [55], and moves to the $k$-space to explain the procedure akin to Wilson’s formulation [56]. In this context, the first peak of the specific heat, $C$, or a marked change in its slope, identifies the most convenient resolution scales where to perform the coarse-graining over the informationally connected nodes, as it allows the iteration of the procedure without an excessive network reduction, and, when the network is scale-invariant, maintaining the dilation symmetry across the renormalization group procedure [37].
@@
To conclude, we can intuitively use the diffusion parameter, $\tau$, to resolve the network topological structures at different resolution scales. Thus, the necessity of analyzing complex networked systems at all levels is solved by resorting to the conjugate Fourier space, the k-space, defined by the Laplacian spectrum and the “wavelengths” of its modes.
II.1.2 Thermodynamics of networks
@@
The statistical physics theory for networks introduced above allows us to view network phenomena from a new perspective and possibly to bridge the gap to network macroscopic observables. In this regard, network thermodynamic quantities such as work, $W$, and heat, $Q$, have been defined in mathematical analogy to their thermodynamic counterparts [1]. Since a network of isolated nodes is the only network with maximum entropy $S_{iso}=\log N$ for $\tau>0$, progressively adding edges diminishes the network’s entropy. Lower entropy reflects the diminished ability of the network to express diversity to a stochastic perturbation, or in other terms, the loss in the internal representation capabilities encoded in the network topological traits. Thus, the entropic cost is given by:
@@
where $0\leq\eta\leq 1$ and $U\equiv-\partial_{\tau}\log Z\equiv\operatorname{Tr}[\hat{L}\hat{\rho}]$. The variational principle here exposed has been devised to macroscopically explain the sparsity and its empirical scaling law, as well as the emergence of topological features such as modularity, small-worldness, and, with partially less clear results, heterogeneity [1]. In our work, we use this variational principle to explain the structural properties of the network subclass composed of heterogeneous hierarchical trees. The specific heat is used to identify the resolution scale at which such a variational principle is effective for our subclass of networks during the transition to the scale-invariant regime. For scale-invariant networks, for some $\tau$-scales, $U=<\lambda>_{\tau}=C_{0}/\tau$ [35, 36], hence in this regime Equation 5 becomes $\eta=\frac{1}{\tau}\frac{C_{0}}{W}=\frac{1}{\tau}\frac{d_{s}}{2W}$. An additional element of this work is that the investigated trees are required to encode the diversity expressed in the environment they are in contact with, defined by a real-world dataset of patterns, $\mathcal{D}$.
II.2 Portegys algorithm
II.2.1 Network growth process
To generate structures based on real-world disorder, we use the algorithm introduced by Portegys [27] for accelerating nearest-neighbor (1-NN) pattern recognition tasks and subsequently extended to k-NN [28]. The Portegys algorithm we use is the extended version with $k=5$, where the distance between patterns is the Euclidean distance.
Given the external constraints, the single control parameter, $r$, and a high-dimensional vector collection, the dataset $\mathcal{D}$, the system seeks the best-fitness configuration to map the dataset’s disorder into a tree, i.e. a network that is loop-less and sparse, $|E|=N-1$, where $|E|$ is the number of links, and $N$ the number of nodes.
The tree-formation protocol reorganizes raw data using a recursive method in which samples are randomly selected and inserted into the growing tree network. At each step, the new pattern is linked to either the current parent pattern or to one of the k-NN children of the parent pattern. The choice is based on the distance between the parent and the new pattern and the one evaluated between the child pattern and the new pattern. The parameter $r$ modulates the relative importance of these two distances creating a larger or smaller region of influence of the parent in the sample space.